Nice rain storm. Collected a full bottle of misty rain overnight and then another full bottle during the day on 9/18. Thick mist associated with the storm.
Monthly Archives: September 2014
20140918 Fritzche
Updates on a few samples and the close down:
9/12/14 11:00 am : full sample present, yet very dirty. Considering how dirty the entire system was, I am unclear if it had been cleaned the previous day by one of my students. Aliquot also taken. Sample was acidified on 9/13.
9/13/14 very small amount – only enough for an aliquot. System cleaned and a blank taken. Blank was acidified that day.
9/14/14 11:15 AM 50 ml of sample – bugs were removed from water and it was acidified that day.
9/18/14 2:00 pm about 100 ml present. Not sure when it may have been collected, since machine had been open for 3 days. Aliquot taken and system powered down. Sample acidified 9/18.
140914 UCSC
Obtained about 100 mL of fog water on Sunday morning at 8:45 am. The direct sunlight had not hit the sample bottle yet.
140913 UCSC
250 mL of clear fog water obtained last night. Split and preserved with HCl.
140910 UCSC
Got about 20 mL last night. collector was open from 9/8 at 9 pm to 9/10 at 11 am when I retrieved the sample. Pretty cloudy water.
140904 UCSC
125 mL of fog water collected. Drizzle and heavy fog
140903 UCSC
125 mL of fog water collected. Drizzle and heavy fog
140829 Pepperwood Observations
140827 LML Observations
Isaac arrived at the collector at 9:30am. The doors to the collector had been deployed however, no sample had been collected. The collector was then disassembled for cleaning. This was preformed by utilizing 409 and DI water. Upon completion of this process the collector was then reassembled and reset for the next fog event.
140830 LML Observations
Isaac arrived at LML at 10:20 am. The doors to the collector had been deployed and a sample had been collected in the sample jar. The sample measured approximately 50 ml. The sample was relatively clear with a small amount of sediment at the bottom. The sample was then removed from the collector and placed in a small labeled bag. The collector was then dissembled and cleaned with 409 and DI water. After the collector was reassembled a blank was then taken which measured approximately 200ml. The collector was then reset for the next fog event.